# Coward: Toward Practical Proactive Federated Backdoor Defense via Collision-based Watermark

### v1

@@ Abstract
Backdoor detection is currently the mainstream defense against backdoor attacks in federated learning (FL), where malicious clients upload poisoned updates that compromise the global model and undermine the reliability of FL deployments. Existing backdoor detection techniques fall into two categories, including passive and proactive ones, depending on whether the server proactively modifies the global model. However, both have inherent limitations in practice: passive defenses are vulnerable to common non-i.i.d. data distributions and random participation of FL clients, whereas current proactive defenses suffer inevitable out-of-distribution (OOD) bias because they rely on backdoor co-existence effects. To address these issues, we introduce a new proactive defense, dubbed Coward, inspired by our discovery of multi-backdoor collision effects, in which consecutively planted, distinct backdoors significantly suppress earlier ones. In general, we detect attackers by evaluating whether the server-injected, conflicting global watermark is erased during local training rather than retained. Our method preserves the advantages of proactive defenses in handling data heterogeneity ($i.e.$, non-i.i.d. data) while mitigating the adverse impact of OOD bias through a revised detection mechanism. Extensive experiments on benchmark datasets confirm the effectiveness of Coward and its resilience to potential adaptive attacks. The code for our method would be available at https://github.com/still2009/cowardFL.
@@
Backdoor Defense, Backdoor Attack, Federated Learning, AI Security, Trustworthy ML.
I Introduction
Federated Learning (FL)  [1, 2, 3] has emerged as a powerful paradigm for privacy-preserving distributed machine learning, enabling collaborative model training among many client devices without sharing raw data. Such character enables secure utilization of massive valuable data from sensitive entities and individuals, unleashing the potential of private data towards personalized and foundational intelligence [4, 5, 6]. It has been widely studied across diverse fields such as healthcare, finance, and edge computing [7, 8, 9].
However, the privacy-preserving nature of federated learning is a double-edged sword. While it safeguards sensitive client data by keeping it local, it also restricts the server’s visibility into client-side behavior. This blind spot opens the door to backdoor attacks [10, 11, 12], which plant malicious pattern into a model such that it behaves normally on clean inputs but produces attacker-desired outcomes when exposed to inputs containing predefined trigger patterns. Exploiting the server’s inability to observe client behavior, malicious participants can covertly implant backdoors into their local models. These compromised updates are then aggregated into the global model, seamlessly propagating the backdoor behavior and undermining the reliability of FL [13, 14, 15].
Figure 1: Comparison between our collision-based proactive defense and the existing method via co-existent mechanism. By developing a backdoor-collided watermark, our method adopts an inverted detection paradigm with respect to malicious clients that is naturally compatible with OOD prediction bias, thereby enabling more reliable and accurate detection.
@@ I Introduction
To tackle this issue, backdoor detection methods has been proposed to identify and exclude malicious updates during model aggregation [16, 17, 18]. Based on the assumption that backdoor training introduces distinct patterns in model parameter space, they achieve detection by identifying such anomalies. Specifically, current techniques mainly use distance-based metrics to assess similarities among client updates and may incorporate clustering algorithms to further reveal structural patterns. In either case, updates identified as outliers are flagged as malicious. However, we reveal that these defense methods are vulnerable to clients’ non-i.i.d. data distribution and participation randomness, which are common settings in practical FL scenarios [3]. These factors lead to highly dynamic and divergent benign updates that may appear as outliers, significantly interfering with detection and leading to a high false positive rate. We refer to this limitation as the non-i.i.d. blurring challenge, as analyzed in Section III-A.
Recently, a new type of defense termed proactive detection, was introduced by the pioneering work, $i.e.$, BackdoorIndicator [18]. Unlike prior works passively rely on client model response, proactive defense actively modifies the global model to provoke divergent behaviors from malicious and benign clients. For instance, BackdoorIndicator injects specific OOD mappings into the global model, which tend to be preserved by malicious clients while being forget by benign clients. As such, the attacker can be identified by inspecting the presence of the injected pattern in local models, where higher inspection accuracy indicates malicious client while lower accuracy indicates benign. By shifting the detection basis from model discrepancies to differences in client behavioral responses, proactive methods naturally avoid the sensitivity to non-i.i.d. data that often undermines traditional defenses.
However, since FL prohibits server access to client training data, current proactive detection methods including Indicator inevitably rely on OOD data in practice, either for injecting patterns during training or evaluating pattern reactions during inference. Yet, deep neural networks (DNNs) are known to produce overconfident and biased predictions on OOD inputs [19, 20, 21], which may undermine the reliability of such detection strategies. Specifically, we reveal that such learning-based proactive methods are fundamentally vulnerable to OOD prediction bias, which causes benign local models to exhibit high inspection accuracy and be misidentified as malicious clients, leading to a high false positive rate. We provide a detailed analysis in Section III-B and refer to this limitation as the OOD bias challenge. Naturally, this raises an important question: can we design a practical proactive detection method that alleviate this limitation?
@@
Fortunately, the answer is yes, although the solution is non-trivial. In this paper, we propose a simple yet effective Collision-based OOD WAtermark for Robust federated backdoor Defense, termed Coward. Our method is inspired by the finding of multi-backdoor collision effect where a consecutively planted different backdoor would significantly suppress the one previously planted. In general, we detect attackers by evaluating whether the server-injected conflicting global watermark has been erased during local training. Clients whose watermark accuracy falls below a specified threshold are then marked as malicious. Specifically, we implant a backdoor-based OOD watermark as the conflicting global watermark, where OOD samples embedded with a trigger are redirected from their original class to a designated watermark target class. We refer to the original class mapping as the base OOD mapping, and the triggered mapping as the watermark mapping. Concretely, the OOD watermark is injected by jointly learning: (1) a regulated base OOD mapping that specifies a fixed class-wise correspondence from OOD samples to the task label space; and (2) a targeted watermark mapping that enforces an association between all triggered OOD samples and the watermark target class. In general, our method alleviates the OOD bias challenge through three aspects: (1) Unlike prior approaches, our collision-based detection treats high inspection accuracy as the signal of benign behavior, making it naturally compatible with the OOD bias–induced high confidence predictions; (2) Our watermark-based detection inspects the presence of the trigger-to-target mapping rather than the OOD mapping, thereby mitigating the direct impact of OOD prediction bias. Together, these designs lead to more reliable backdoor identification and lower false positive rates. and (3) our base OOD mapping can mitigates OOD prediction bias by enforcing explicit class-wise assignments to OOD samples; We demonstrate (1) and (2) in Figure1, and verify (3) in Figure7.
@@
In summary, our contributions are three-fold: (1) We revisit the limitations of existing methods by identifying and analyzing two key challenges unique to federated learning: non-i.i.d. blurring challenge and OOD prediction bias. (2) We find an intriguing phenomenon, dubbed the multi-backdoor collision effect, where a consecutively planted different backdoor would significantly suppress the previous implanted one. Based on this finding, we design a simple yet effective proactive federated backdoor defense method, $i.e.$, Coward, that effectively mitigates OOD bias and achieves more accurate detection. (3) Extensive experiments on benchmark datasets verify and explain the effectiveness of our Coward method and its resistance to potential adaptive attacks.
II Background and Related Works
II-A Federated Learning
@@
Federated learning (FL) enables multiple decentralized clients to collaboratively train a shared model without exchanging their private data [2]. Based on how data is partitioned, FL is typically categorized into horizontal FL (HFL), where clients share the same feature space but different samples, and vertical FL (VFL), where clients share samples but differ in input feature fields [22].
@@
In this paper, we focus on the HFL setting, where a central server coordinates model training and aggregation in iterative rounds. In each round, a subset of clients is selected to receive the current global model, train it on their local data, and send back the updated parameters. The server then aggregates these updates to form a new global model. This process repeats, gradually improving model performance while keeping data decentralized. Specifically, given the selected client set $\mathcal{S}^{t-1}$ in round $t-1$, the global model $\bm{w}^{t}$ in round $t$ is:
@@
where $N_{k}$ and $N$ denote the data volume of client $k$ and the total data volume across participating clients, respectively. $\Delta$ denotes the model residual with respect to the previously broadcasted global model. We highlight three key characteristics that distinguish FL from traditional centralized settings:
@@
  • Non-i.i.d. Client Data [1]: Data distributions vary significantly across clients, resulting in inconsistent local updates and increased optimization difficulty.
@@
  • Partial Client Participation [1]: In each training round, only a subset of clients is selected due to limited communication resources and dynamic client availability.
These characteristics introduce unique challenges for both attacks and defenses in FL, as described in subsequent parts.
II-B Backdoor Attacks in Federated Learning
@@
Backdoor attacks aim to implant hidden behaviors into neural networks, causing them to function normally on clean data but misclassify inputs containing attacker-specified triggers [10, 23]. Data poisoning is the most common backdoor injection method, where a subset of training samples is embedded with triggers and mislabeled to a target class [23]. Stealth and persistence are key objectives in backdoor attacks, motivating the design of invisible [24, 25], semantic [26, 27], optimized [28, 29] and sample-specific [30, 31, 32] triggers. However, these approaches are limited when directly applied to FL, due to the unique challenge that adversaries can only manipulate a limited subset of clients and must ensure their backdoored updates survive the aggregation process to impact the global model. Thus, FL-specific backdoor attacks have emerged, broadly classified into following two categories.
@@
Single-client Attacks. The first backdoor attack for FL introduced the model replacement attack [13], where a malicious client uploaded a scaled local update to dominate the aggregation. However, the abnormally large update made it easily detectable under defense [33]. To improve stealth and persistence, follow-up works introduced more subtle attack strategies. For example, PGD [33] constrained malicious updates within an $\ell_{2}$-ball to evade norm-based defenses, while Neurotoxin [34] injected backdoors into parameters associated with inactive neurons to minimize noticeable changes. Other works aimed to further blend backdoor patterns into normal behavior. For example, Edge-case attacks [35] utilized naturally occurring features as triggers, reducing the likelihood of detection, while Chameleon [11] aligned backdoor and benign features in latent space using contrastive learning.
@@
Multi-client Attacks. The DBA framework [36] leveraged the distributed nature of FL to spread the backdoor across multiple clients, with each client using a distinct local trigger. These independently subtle triggers collectively formed an effective global backdoor, reducing the detectability of any individual client. Expanding on this idea, Gong et al. [37] proposed a coordinated multi-client attack, in which each local trigger was optimized in a model-aware fashion. This coordination allowed the attackers to craft triggers that were both less suspicious and more effective, further amplifying the threat posed by distributed poisoning. In contrast, Non-Cooperative Backdoor Attacks (NBA) [38] involved independent attackers using distinct triggers for different targets, yet still exhibited enhanced stealth comparable to coordinated methods.
II-C Backdoor Defenses in Federated Learning
@@
In centralized settings, backdoor defenses typically focus on either removing the backdoor from the model [39, 40] or filtering trigger samples during training or inference [41, 42]. However, these approaches often become impractical in FL due to the limited capabilities of the defender ($i.e.$, the central server). Specifically, (1) the server lacks access to clean and in-distribution data, making data-dependent removal methods like pruning, retraining, or fine-tuning infeasible. (2) the server cannot control the local training process, rendering sample filtering and robust training strategies infeasible. To address these limitations, FL-specific defense mechanisms have been developed, generally falling into below two categories.
@@
Backdoor Effect Mitigation methods aim to reduce the impact of backdoors in the global model by suppressing anomalous neurons across local updates. Early works such as Weak-DP [33] applied norm clipping and noise injection to weaken abnormal neuron effects. Statistical approaches like geometric median [43] and Trimmed-Mean [44] enforced robustness by aggregating updates in a statistically neutral manner. RLR [45] further enhanced mitigation granularity by assigning parameter-wise adaptive learning rates. FLPurifier [46] achieved stronger backdoor suppression by performing adaptive aggregation at the classifier level under a modified FL protocol that decoupled encoder and classifier training. However, these methods operate all client updates equally, often diminishing benign contributions and harming overall model performance under the standard FL protocol. This limitation has led to the rise of the following detection-based approaches as a mainstream alternative.
[…]
Building on the shared characteristic that prior methods rely on passive server-side detection without any preparatory intervention, we categorize them as passive detection. In contrast, proactive detection, such as Indicator [18] and our proposed approach, take an active role by modifying the global model to provoke distinguishable behaviors from clients. By leveraging multi-backdoor interaction mechanisms that is less sensitive to data heterogeneity, these methods improve detection robustness. While Indicator marks a key advance and outperforms passive baselines, as we will show in the next section, it still suffers from OOD bias and high false positives. How to design an effective proactive federated backdoor defense in practice is still an important open question.
@@
III Revisiting Existing Defenses and Multi-Backdoor Interactions
In this section, we first revisit the limitations of existing passive and proactive defenses, respectively. We then delve into the multi-backdoor interaction effects, a key underlying mechanism that drives the design of proactive defenses.
III-A Revisiting Passive Defenses

### v2

# Collision-based Watermark for Detecting Backdoor Manipulation in Federated Learning

@@ Abstract
As AI-generated content increasingly underpins real-world applications, its accompanying security risks, including privacy leakage and copyright infringement, have become growing concerns. In this context, Federated Learning (FL) offers a promising foundation for enhancing trustworthiness by enabling privacy-preserving collaborative learning over proprietary data. However, its practical adoption is critically threatened by backdoor-based model manipulation, where a small number of malicious clients can compromise the system and induce harmful content generation and decision-making. Although various detection methods have been proposed to detect such manipulation, we reveal that they are either disrupted by non-i.i.d. data distributions and random client participation, or misled by out-of-distribution (OOD) prediction bias, both of which are unique challenges in FL scenarios. To address these issues, we introduce a novel proactive detection method dubbed Coward, inspired by our discovery of multi-backdoor collision effects, in which consecutively planted, distinct backdoors significantly suppress earlier ones. Correspondingly, we modify the federated global model by injecting a carefully designed backdoor-collided watermark, implemented via regulated dual-mapping learning on OOD data. This design not only enables an inverted detection paradigm compared to existing proactive methods, thereby naturally counteracting the adverse impact of OOD prediction bias, but also introduces a low-disruptive training intervention that inherently limits the strength of OOD bias, leading to significantly fewer misjudgments. Extensive experiments on benchmark datasets show that Coward achieves state-of-the-art detection performance, effectively alleviates OOD prediction bias, and remains robust against potential adaptive manipulations. The code for our method is available at https://github.com/still2009/cowardFL.
@@
Backdoor Detection, Backdoor Manipulation, Federated Learning, Trustworthy Machine Learning, AI Security.
1 Introduction
AI-generated content, including text, images, numerical predictions, and categorical decisions, has become ubiquitous and now serves as a core component of modern intelligent applications [cao2025survey]. However, growing risks of harmful content generation [shen2025hatebench, ba2024exposing], malicious manipulation [chu2025jailbreakradar, shensynthetic], data privacy abuse [carlini2021extracting, chen2025survey], and copyright infringement [reuters2023_openai_copyright, liu2025harnessing], highlight an urgent need to develop trustworthy AI content generation and decision systems. As a promising starting point, Federated Learning (FL) [mcmahan2017fedavg, fl_survey24_pami] provides a trustworthy collaborative learning paradigm that enables the privacy-preserved utilization of cross-organization domain data. By keeping raw data locally owned and processed, FL naturally strengthens data privacy and copyright protection, while also unlocking the value of proprietary, professional, and domain-specific data that are inaccessible to models trained solely on generic public knowledge. Together, these properties position FL as an increasingly attractive solution for building trustworthy foundation models [fl_fm_nips, fan2025_fl_foudationmodel], and have motivated its wide exploration across various domains such as healthcare, finance, and edge intelligence [fl_nature_medi24, fl_finance_icse24, fl_edge_survey24].
Figure 1: Our backdoor-collided detection versus the existing backdoor-coexistent method. By developing a backdoor-collided watermark, our method enables an inverted detection paradigm that remains effective under OOD prediction bias, a key challenge that limits current proactive detection methods.
However, the privacy-preserving design of federated learning is a double-edged sword. While keeping raw data local effectively protects sensitive information, it also restricts the server’s ability to inspect participant behavior during training. This restricted observability creates a structural vulnerability to model-level manipulation, where the integrity of the federated global model can be compromised by covertly manipulated local models, most prominently through backdoor attacks [ba_survey_yiming, chameleon]. In general, backdoor attacks manipulate models to behave normally on benign inputs while producing harmful outputs only when specific trigger conditions are met, making them more stealthy and difficult to perceive. Once the manipulated local models are aggregated together with benign ones, the backdoor is seamlessly propagated to the global model, thereby manipulating the entire federated system [bagdasaryan2020backdoor, tifs25_fl_atk2].
@@ 1 Introduction
To detect and exclude backdoor-manipulated models, passive detection methods have been proposed [rieger2022deepsight, nguyen2022flame], which assume that malicious models expose observable model-level deviations as outliers. They rely on similarity-based measurements over either model parameters or model outputs and identify anomalous client models either through direct thresholding or by further applying clustering to expose structurally isolated outliers. However, we reveal that these methods are vulnerable to clients’ non-i.i.d. data distribution and participation randomness, which are common settings in FL scenarios [fl_survey24_pami]. These factors lead to highly dynamic and divergent benign local models that may appear as outliers, significantly interfering with detection. We refer to this limitation as the non-i.i.d. blurring challenge, as analyzed in Section 3.2.
Recently, the pioneering work BackdoorIndicator [li2024backdoorindicator] proposed the first federated proactive detection strategy, which explicitly intervenes in the training process to induce manipulated models to exhibit behaviors distinctive from benign ones. Specifically, BackdoorIndicator injects OOD patterns specifically designed to be easily retained by manipulated clients while tending to be forgotten by benign ones, such that stronger pattern retention in a local model indicates a manipulated client. By shifting the detection basis from model discrepancies to differences in client responses to intervention actions, proactive methods naturally avoid the sensitivity to non-i.i.d. data that often undermines traditional defenses. However, due to the privacy constraints of FL that prevent access to in-distribution client data, proactive training interventions at the server are restricted to perform via OOD data. As a result, detection outcomes become susceptible to the OOD prediction bias of deep neural networks (DNNs) [nguyen2015ood_bias, guo2017ood_bias_calib, hein2019relu_ood_bias_theory], causing benign local models to be easily misidentified as malicious clients. We provide a detailed analysis in Section 3.3 and refer to this limitation as the OOD bias challenge.
@@
To tackle above challenges, we propose a simple yet effective Collision-based OOD WAtermark for Robust federated backdoor Detecton method, termed Coward. Our method is inspired by the finding of multiple backdoor manipulation collision effect where a consecutively planted different backdoor would significantly collide the one previously planted. To this end, we detect model manipulation by evaluating whether the server-injected conflicting global watermark has been erased during local training. Specifically, we implant a conflicting OOD watermark, where OOD samples embedded with a specific watermark trigger are redirected from their original class to a designated target class. Concretely, the OOD watermark is injected by jointly learning: (1) a regulated base OOD mapping that specifies a fixed class-wise correspondence from OOD samples to the task label space; and (2) a targeted watermark mapping that enforces an association between all triggered OOD samples and the watermark target class. In general, our method alleviates the OOD bias challenge through three aspects: (1) Unlike prior approaches, our collision-based detection treats high inspection accuracy as the signal of benign behavior, making it naturally compatible with the OOD bias–induced high confidence predictions; (2) Our watermark-based detection inspects the presence of the trigger-to-target mapping rather than the OOD sample-to-target mapping, mitigating the direct impact of OOD prediction bias. and (3) our method enforces explicit class-wise label assignments to OOD samples to reduce random prediction fluctuations; We demonstrate (1) and (2) in Figure1, and verify (3) in Figure7.
In summary, our contributions are three-fold:
  • We revisit existing detection methods against FL backdoor manipulation and identified two key challenges limiting their effectiveness: non-i.i.d. blurring challenge and OOD prediction bias unique characterized under FL sceanrios.
@@
  • We find an intriguing phenomenon, dubbed the multi-backdoor manipulation collision effect, where a consecutively planted different backdoor would significantly suppress the previous implanted one. Based on this finding, we design a simple yet effective proactive backdoor manipulation detection method, $i.e.$, Coward, that effectively mitigates OOD bias and achieves more accurate detection.
  • Extensive experiments on benchmark datasets verify the effectiveness of our method and its resistance to potential adaptive backdoor attacks.
2 Background and Related Works
2.1 Federated Learning
@@
Federated learning enables multiple decentralized clients to collaboratively train a shared model without exchanging their private data [fl_survey24_pami]. Based on how data is partitioned, FL is typically categorized into horizontal FL (HFL), where clients share the same feature space but different samples, and vertical FL (VFL), where clients share samples but differ in input feature fields [flsurvey_yangqiang].
@@
In this paper, we focus on the HFL setting, where a central server coordinates model training by periodically selecting a subset of clients to receive the current global model, update it using local data, and return the updated parameters. The server then aggregates these updates to form a new global model, and this process repeats iteratively, progressively improving model performance while keeping all training data local. Specifically, given the selected client set $\mathcal{S}^{t-1}$ in round t1t-1, the global model $\bm{w}^{t}$ in round $t$ is:
@@
where $N_{k}$ and $N$ denote the data volume of client $k$ and the total data volume across participating clients, respectively. $\Delta$ denotes the model residual with respect to the previous global model. We highlight three key characteristics that distinguish FL from centralized settings:
@@
  • Non-i.i.d. Client Data [mcmahan2017fedavg]: Data distributions vary significantly across clients, resulting in inconsistent local updates and increased optimization difficulty.
@@
  • Partial Client Participation [mcmahan2017fedavg]: In each training round, only a subset of clients is selected due to limited communication resources and dynamic client availability.
2.2 Backdoor Manipulation in Federated Learning
@@
Backdoor attacks aim to implant hidden behaviors into neural networks, causing them to function normally on clean data but misclassify inputs containing attacker-specified triggers [ba_survey_yiming, gu2020badnets]. Data poisoning is the most common backdoor injection method, where a subset of training samples is embedded with triggers and mislabeled to a target class [gu2020badnets]. Stealth and persistence are key objectives in backdoor attacks, motivating the design of invisible [nguyen2021wanet, Gao_attack24], semantic [bagdasaryan2021blind, xu2024csba_semantic_atk], optimized [geiping2020witches, sun2024backdoor] and sample-specific [iad, li2021invisible, zhu2025towards] triggers. However, these approaches are limited in FL, because adversaries can only manipulate a limited subset of clients and must ensure their backdoored updates survive the aggregation process to impact the global model. Specifically, FL-specific backdoor manipulation can be classified into following two categories.
@@
Single-client Manipulation. The first backdoor manipulations method for FL introduced the model replacement attack technic  [bagdasaryan2020backdoor], where a malicious client uploaded a scaled local update to dominate the aggregated model. However, the abnormally large update made it easily detectable under detection [zitengsun01pgd]. To improve stealth and persistence, follow-up works introduced more subtle manipulation strategies. For example, PGD [zitengsun01pgd] constrained malicious updates within an $\ell_{2}$-ball to evade norm-based detection, while Neurotoxin [Zhengming01Neurotoxin] injected backdoors into parameters associated with inactive neurons to minimize noticeable changes. Other works aimed to further blend backdoor patterns into normal behavior. For example, Edge-case manipulation [wang2020edgecase] utilized naturally occurring features as triggers, reducing the likelihood of detection, while Chameleon [chameleon] aligned backdoor and benign features using contrastive learning.
@@
Multi-client Manipulation. The DBA framework [xie2019dba] leveraged the distributed nature of FL to spread the backdoor across multiple clients, with each client using a distinct local trigger. These independently subtle triggers collectively formed an effective global backdoor, reducing the detectability of any individual client. Expanding on this idea, Gong et al. [gong2022coordinated_dba] proposed a coordinated multi-client manipulation method, in which each local trigger was optimized in a model-aware fashion. This coordination allowed the attackers to craft triggers that were both less suspicious and more effective, further amplifying the threat posed by distributed poisoning. In contrast, Non-Cooperative Backdoor Attacks (NBA) [nguyen2024nba] involved independent attackers using distinct triggers for different targets, yet still exhibited enhanced stealth comparable to coordinated methods.
2.3 Backdoor Detection in Federated Learning
@@
In centralized settings, backdoor defense typically focus on either removing the backdoor from the model [li2021ba_defense_kd, wu2021ba_defense_prune] or filtering trigger samples during training or inference [qiu2021deepsweep, javaheripi2020cleann]. However, these approaches are often impractical in FL due to the defender’s limited capabilities ($i.e.$, the central server): (1) the server lacks access to clean and in-distribution data, making data-dependent removal methods like pruning, retraining, or fine-tuning infeasible. (2) the server cannot control the local training process, rendering sample filtering and robust training strategies infeasible. To address these limitations, FL-specific backdoor detection mechanisms have been developed, generally falling into below two categories.
@@
Backdoor Effect Mitigation methods aim to reduce the impact of backdoor behaviors in the global model by suppressing anomalous neurons across local models. Early works such as Weak-DP [zitengsun01pgd] applied norm clipping and noise injection to weaken abnormal neuron effects. Statistical approaches like geometric median [RFA] and Trimmed-Mean [blanchard2017machine] enforced robustness by aggregating updates in a statistically neutral manner. RLR [ozdayi2021RLR] further enhanced mitigation granularity by assigning parameter-wise adaptive learning rates. FLPurifier [zhang2024flpurifier] achieved stronger backdoor suppression by performing adaptive aggregation at the classifier level under a modified FL protocol that decoupled encoder and classifier training. However, these methods operate all clients equally, often diminishing benign model contributions and harming overall performance under the standard FL protocol. This limitation has led to the rise of the following detection-based methods as a mainstream alternative.
Backdoor Client Detection methods aim to identify and exclude manipulated local models during aggregation by quantifying the differences between benign and malicious clients, typically through distance-based metrics on model parameters. Clients whose update patterns deviate significantly from the majority are flagged as malicious. For instance, Multi-Krum [multiKrum] leveraged Euclidean distance, while Foolsgold [foolsgold], FLTrust [cao2021fltrust], and ShieldFL [ma2022shieldfl] adopted cosine similarity, all applied directly on full model weights. Rflbat [wang2022rflbat] extended this by computing distances in a top-$k$ PCA-projected space, aiming to reduce noise and improve efficiency, though at the cost of some information loss. Beyond model parameters, DeepSight [rieger2022deepsight] incorporated prediction-level discrepancies of local models, enabling a more comprehensive analysis. Extending beyond pure detection, Flame [nguyen2022flame] combined clustering-based client identification with gradient norm-based mitigation, positioning itself at the intersection of detection and mitigation strategies and demonstrating strong empirical performance. FLGuardian [tifs25_fl_defens_poison], a recent study addressing untagged backdoor manipulation, performed layer-wise model analysis using similarity metrics and clustering algorithms, reflecting a growing shift toward more fine-grained detection in the parameter space. However, these methods assume malicious models are outliers, making them sensitive to data heterogeneity and ineffective against stealthy attacks. Since these detection methods operate without any server-side intervention, we categorize them as passive detection. In contrast, proactive detection explicitly intervenes in FL training to facilitate backdoor identification. The recent pioneering work Indicator [li2024backdoorindicator] adopts this strategy by injecting OOD mappings into the global model, achieving improved robustness under data heterogeneity. However, it still suffers from high false-positive rates due to OOD prediction bias, leaving the design of effective proactive federated backdoor detection methods an open challenge.
@@
3 Revisiting Existing Detection Methods and Multi-Backdoor Manipulation Interplay
In this section, we first introduce the threat model and revisit the limitations of existing detection methods. We then examine the interplay effects of multi-backdoor manipulation, which directly motivate our detection method.
3.1 Preliminary: Threat Model
Attacker’s Goal and Capability. The attacker resides on the client side and aims to manipulate the federated system by injecting a targeted backdoor into the global model, causing it to misclassify all trigger-embedded inputs into a designated target class while maintaining normal performance on benign inputs. Although the attacker cannot directly access the global model, they may control a subset of clients and submit arbitrary malicious models or gradients to the server, thereby influencing the global model updates.
3.2 Revisiting Passive Detection